OpenStudy (candi5a18):

How do you determine horizontal,vertical and slant asymptotes for rational functions?

OpenStudy (biue):

?

Hayhayz (hayhayz):

Hi Welcome to Openstudy!

OpenStudy (irishboy123):

a function is a rational function if looks like this: \[ f(x) = \frac{P(x)}{Q(x)} \] a *ratio* of functions... kinda obvious, maybe, but does that help?!

OpenStudy (irishboy123):

so think: Q = 0 or P = 0 or P = Q

OpenStudy (candi5a18):

What if P=Q

OpenStudy (comrad):

Well an asymptote is where the graph of the equation does not touch

OpenStudy (comrad):

So if a graph doesn't touch the x axis, say the graph is a parabola. Your asymptote would be y=x

OpenStudy (irishboy123):

you have \(\dfrac{0}{0}\) and that's where you earn your money.....!! but try to think of an example of that ......eg \(\dfrac{sin (x) }{x}\)...... and there is often a solution.

OpenStudy (candi5a18):

I have a general idea on how to get horizontal and vertical asymtotes, I am confused about slant ones

OpenStudy (candi5a18):

\[f(x)=\frac{x^2+2x+1}{x+2}\] What is the slant asymptote for this function?

OpenStudy (irishboy123):

 like that?!